Twin sisters, Sue and Debbie, are both good math students. their mathematics teacher calculated that sue has a probability of 0.7 of getting an A at the end of the year and Debbie has a probability of 0.6. Find the probability that:
i) Neither gets an A.
ii) if only one gets an A , its is Sue.

Answer by stanbon(75887)

Twin sisters, Sue and Debbie, are both good math students. their mathematics teacher calculated that sue has a probability of 0.7 of getting an A at the end of the year and Debbie has a probability of 0.6. Find the probability that:
i) P(Neither gets an A) = (1-0.7)(1-0.6) = 0.3*0.4 = 0.12
---------------------------
ii) if only one gets an A , it is Sue.
P(both get an A) = 0.7*0.6 = 0.42
-------------------------------------
P(one gets an A) = = 0.7+0.6-0.42 = 1.3-0.42 = 0.88
-------------------------------------
P(only one gets A) = 1-[0.12+0.42] = 1-0.54 = 0.46
---------------------------------------------------
P(Sue gets A | one gets A) = P(one gets A AND Sue gets A)/P(one gets A)
---
= P(Sue gets A)/P(one gets A) = 0.7/0.88 = 0.7955
=================
Cheers,
Stan H.